# Milliarcsecond astrometric oscillations in AGN as a precursor of multi-messaging gravitational wave events

### v1

@@ 6 Summary and conclusions
Astrometric VLBI monitoring with sufficient angular resolution holds significant potential for advancing the study of SMBHBs during the hardening phase of their evolutionary pathway, and possibly even during the immediate pre-coalescence stage. The brief and by no means exhaustive overview of the SMBHB landscape provided in Section 1 underscores the importance of each new SMBHB detection. Particularly valuable are those candidates identified in the late stages of their evolution, approaching coalescence, as these represent prime targets for multi-messenger observations. Although astrometric oscillations in active galactic nuclei (AGN) cannot, in isolation, serve as conclusive evidence for the presence of an SMBHB, such systems remain among the most plausible explanations for the observed astrometric behaviour. We also briefly outline the prospect of directly imaging SMBHBs in advanced evolutionary stages as a compelling scientific objective for future space-based mm/sub-mm VLBI missions. We anticipate that the forthcoming detection of gravitational waves by LISA, when combined with astrometric VLBI monitoring, will significantly advance our ability to confirm the existence and characterise the properties of SMBHBs.
@@
The study presented in this paper is based on ad hoc peruse of VLBI astrometric measurements conducted for purposes other than search for possible SMBHB systems. The major feature attracted our attention was the apparent regular oscillations of the sky position of the sources J21026015 (reported in our previous paper Titov et al. (2023)) and the new case of J02041514 presented in the current work) with sub-millarcsecond amplitudes. We focused our attention on the latter source as a case study and interpret the oscillating astrometric pattern as a manifestation of orbital motion in SMBHB system with the period of yr in the observer’s frame. We underline that none of these two sources is confirmed beyond doubts as containing SMBHBs. Nevertheless, a Keplerian motion of components in the suspected SMBHBs is a very plausible explanation of the observed astrometric oscillations. By stating so, we emphasise that our study does not and cannot rule out another reason for the apparent astrometric behaviour, e.g., variations of the distribution of radio brightness in the sources due to their jet dynamics. A more careful re-analysis of massive available astrometric VLBI data for other sources, including several dozens of sources reported by Makarov et al. (2024), might provide indications on new SMBHB candidates.
@@
Prompted by our interest in potential presence of SMBHB in the object J21026015 – whose oscillatory behaviour was presented in our previous study (Titov et al., 2023) – a careful re-examination of the available optical spectrum has led us to propose in Appendix A a more plausible redshift value of , in place of the value previously adopted. This might be useful for future detailed studies of this source.
@@
The method of evaluation of astrometric VLBI time series presented in the current study allowed us to detect a harmonic oscillating astrometric pattern of J02041514, with the apparent period of yr. We emphasise that the goal of the present study was to describe quantitatively how observed astrometric oscillations of AGN containing an SMBHB can be used for evaluating GW generation by such an object. This ‘multi-messaging’ link between astrometric patterns detected in radio domain and GW is of fundamental value and does not depend on specific quantitative values of astrometric oscillations (periods).
@@
Assuming that the detected oscillating astrometric pattern is a manifestation of SMBHB, we develop a ‘toy’ model of such the system which is linked to the observationally detected astrometric oscillations with just one parameter, the oscillations period. The toy model of J02041514 allows us to investigate the evolution of orbital motion of SMBHB and predict its manifestations as a multi-messaging source generating gravitational waves. We quantitatively evaluate potential detectability of generated by our source GW emission by the LISA gravitational wave telescope and contribution of SMBHB with parameters of our toy model into the GW background detected by Pulsar Timing Arrays.
@@ Acknowledgements.
- Astropy Collaboration et al. (2018) Astropy Collaboration, Price-Whelan, A. M., Sipcz, B. M., et al. 2018, The Astronomical Journal, 156, 123
- Astropy Collaboration et al. (2013) Astropy Collaboration, Robitaille, T. P., Tollerud, E. J., et al. 2013, Astronomy & Astrophysics, 558, A33
@@
- Baluev (2008) Baluev, R. V. 2008, Monthly Notices of the Royal Astronomical Society, 385, 1279
@@
- D’Orazio & Charisi (2023) D’Orazio, D. J. & Charisi, M. 2023, arXiv e-prints, arXiv:2310.16896
@@
- Gross et al. (2024) Gross, A. C., Chen, Y.-C., Oguri, M., et al. 2024, arXiv [arXiv:2409.16363]
@@
- Gurvits et al. (2022) Gurvits, L. I., Paragi, Z., Amils, R. I., et al. 2022, Acta Astronautica, 196, 314
- Gurvits et al. (2021) Gurvits, L. I., Paragi, Z., Casasola, V., et al. 2021, Experimental Astronomy, 51, 559
@@
- Jones et al. (2018) Jones, K. M., Ghosh, T., & Salter, C. J. 2018, The Astronomical Journal, 155, 254
@@ Appendix A On the redshift of the source J21026015
In the discovery paper by Sowards-Emmerd et al. (2004), a low dispersion spectrum of the object taken with the 9.2-m Hobby Eberly Telescope was presented. It shows a strong emission line at Å, which has been interpreted as a Ly emission line. However, a more careful examination of the spectrum allows us to conclude that there is no strong Lyman break absorption blueward of the Ly emission line typical of high-redshift quasars. The usually strong C IV 1549 Å line was not detected either. Therefore, the emission line is not likely to be that of a high-redshift Ly. With only a single emission line in the spectrum, the exact redshift cannot be uniquely determined. However, due to the lack of other strong lines in the wavelength range covered by the observed spectrum, most likely this line corresponds to the Mg ii 2800 Å transition. According to this interpretation, the redshift of this source is .

### v2

@@ 6 Summary and conclusions
Astrometric VLBI monitoring with a sufficient angular resolution holds significant potential for advancing the study of SMBHBs during the hardening phase of their evolutionary pathway, and possibly even during the immediate pre-coalescence stage. The brief and by no means exhaustive overview of the SMBHB landscape provided in Section 1 underscores the importance of each new SMBHB detection. Particularly valuable are those candidates identified in the late stages of their evolution, approaching coalescence, as these represent prime targets for multi-messenger observations. Although astrometric oscillations in AGNs cannot, in isolation, serve as conclusive evidence for the presence of an SMBHB, such systems remain among the most plausible explanations for the astrometric behaviour observed. We also briefly outline the prospect of directly imaging SMBHBs in advanced evolutionary stages as a compelling scientific objective for future space-based millimetre/sub-millimetre VLBI missions. We anticipate that the forthcoming detection of GWs by LISA, when combined with astrometric VLBI monitoring, will significantly advance our ability to confirm the existence and characterise the properties of SMBHBs.
@@
The study presented in this paper is based on ad hoc peruse of VLBI astrometric measurements conducted for purposes other than a search for possible SMBHB systems. The major feature that attracted our attention was the apparent regular oscillations of the sky position of the source J21026015 (reported in our previous paper Titov et al. (2023)) and the new case of J02041514 presented in the current work) with sub-millarcsecond amplitudes. We focus our attention on the latter source as a case study and interpret the oscillating astrometric pattern as a manifestation of orbital motion in a SMBHB system with a period of yr in the observer’s frame. We underline that neither of these sources is confirmed beyond doubt as containing SMBHBs. Nevertheless, a Keplerian motion of components in the suspected SMBHBs is a very plausible explanation of the observed astrometric oscillations. By stating this, we emphasise that our study does not and cannot rule out another reason for the apparent astrometric behaviour; for example, variations in the distribution of radio brightness in the sources due to their jet dynamics. A more careful re-analysis of abundant available astrometric VLBI data for other sources, including several dozen sources reported by Makarov et al. (2024), might provide indications of new SMBHB candidates.
@@
Prompted by our interest in the potential presence of a SMBHB in the object J21026015 – whose oscillatory behaviour was presented in our previous study (Titov et al. 2023) – a careful re-examination of the available optical spectrum has led us to propose in Appendix A a more plausible redshift value of , in place of the value previously adopted. This might be useful for future detailed studies of this source.
@@
The method of evaluating astrometric VLBI time series presented in the current study allowed us to detect a harmonic oscillating astrometric pattern for J02041514, with an apparent period of yr. We emphasise that the goal of the present study is to describe quantitatively how observed astrometric oscillations of AGNs containing an SMBHB can be used to evaluate GW generation by such an object. This multi-messenger link between astrometric patterns detected in the radio domain and GWs is of fundamental value and does not depend on specific quantitative values of astrometric oscillations (periods).
@@
Assuming that the detected oscillating astrometric pattern is a manifestation of SMBHBs, we develop a toy model of such a system that is linked to the observationally detected astrometric oscillations with just one parameter, the oscillation period. The toy model of J02041514 allows us to investigate the evolution of orbital motion of a SMBHB and predict its manifestations as a multi-messenger source generating GWs. We quantitatively evaluate the potential detectability of GW emission generated by our source by the LISA GW telescope and the contribution of a SMBHB with the parameters of our toy model to the GW background detected by PTAs.
@@ Acknowledgements.
- Astropy Collaboration et al. (2018) Astropy Collaboration, Price-Whelan, A. M., Sipcz, B. M., et al. 2018, AJ, 156, 123
- Astropy Collaboration et al. (2013) Astropy Collaboration, Robitaille, T. P., Tollerud, E. J., et al. 2013, A&A, 558, A33
@@
- Baluev (2008) Baluev, R. V. 2008, MNRAS, 385, 1279
@@
- D’Orazio et al. (2024) D’Orazio, D. J., Charisi, M., Derdzinski, A., et al. 2024, in Black Holes in the Era of Gravitational-Wave Astronomy, ed. M. Arca Sedda, E. Bortolas, & M. Spera (Elsevier), 379–478, (updated version: arXiv:2310.16896)
@@
- Gross et al. (2024) Gross, A. C., Chen, Y.-C., Oguri, M., et al. 2024, arXiv [arXiv:2409.16363], ApJ, submitted
@@
- Gurvits et al. (2022) Gurvits, L. I., Paragi, Z., Amils, R. I., et al. 2022, Acta Astron., 196, 314
- Gurvits et al. (2021) Gurvits, L. I., Paragi, Z., Casasola, V., et al. 2021, Exp. Astronomy, 51, 559
@@
- Jones et al. (2018) Jones, K. M., Ghosh, T., & Salter, C. J. 2018, AJ, 155, 254
@@ Appendix A On the redshift of the source J21026015
In the discovery paper by Sowards-Emmerd et al. (2004), a low dispersion spectrum of the object taken with the 9.2-m Hobby Eberly Telescope was presented. It shows a strong emission line at Å, which has been interpreted as a Ly emission line. However, a more careful examination of the spectrum allows us to conclude that there is no strong Lyman break absorption bluewards of the Ly emission line typical of high-redshift quasars. The usually strong C IV 1549 Å line was not detected either. Therefore, the emission line is not likely to be that of a high-redshift Ly. With only a single emission line in the spectrum, the exact redshift cannot be uniquely determined. However, due to the lack of other strong lines in the wavelength range covered by the observed spectrum, most likely this line corresponds to the Mg ii 2800 Å transition. According to this interpretation, the redshift of this source is .